OpenStudy (anonymous):

Help! need step by step instructions. Will become a fan and give you a medal (: Your car gets 25 mi/gal around town and gets 30 mi/gal on the highway. a. If 50% of the miles you drive are around town and 50% of the miles you drive are on the highway, what is your overall average miles per gallon? b. If 60% of the miles you drive are on the highway and 40% is around town, what is your average miles per gallon?

OpenStudy (anonymous):

a. 30+25=55 55/2=27.5 27.5 is your average mpg b. 25x4=100, & 30x6=180. 180+100=280 280/10=28 28 is your average mpg